OpenStudy (anonymous):

solve in special products: (m+n)(m-n)(m^2+n^2)(m^4+n^4)(m^8+n^8)(m^16+n^16)

OpenStudy (sriram):

=(m^2-n^2)(m^2+n^2)(m^4+n^4)(m^8+n^8)(m^16+n^16) =(m^4-n^4)(m^4+n^4)(m^8+n^8)(m^16+n^16) =(m^8-n^8)(m^8+n^8)(m^16+n^16) doing so we end up with m^32-n^32

OpenStudy (anonymous):

thanks